How many electrons constitute 1 microC of charge ............  .    ?

Grade:12

1 Answers

Khimraj
5 years ago

You can easily find the answer of this question by using the formula.. Q=ne ; where 'Q’ is the total charge, 'n’ is no. of electrons and 'e’ is the charge of one electron.

Here, Q=10^-6 C and e=1.6 x 10^-19 C.

So, n= Q/e =10^-6/1.6 x 10^-19 = 6.25 x 10^12

So,there are 6.25 x 10^12 electrons present in 1 microcoulomb of charge.